OpenStudy (anonymous):

A train, starting at 11 am, travels east at 45 mph while another, starting at noon from the same point, travels south at 60 mph. How fast are they separating at 3pm?

OpenStudy (mrdoe):

Pythagorean theorem

OpenStudy (anonymous):

how

OpenStudy (mrdoe):

go ahead and plug in your values to find out how fast each is separating in it's direction, use a^2 + b^2 = c^2 to find how fast the distance between them is increasing

OpenStudy (mrdoe):

just think about the trains as vectors

OpenStudy (anonymous):

i got 75

OpenStudy (anonymous):



OpenStudy (anonymous):

show the steps please!

OpenStudy (anonymous):

\[(45+x)^2+y^2=d^2\] \[2(45+x)x'+2yy'=2dd'\] you want \(d'\) and you are told \(x'=45,y'=60\) plug in the numbers, solve for \(d'\)

OpenStudy (anonymous):

why it is 45+x

OpenStudy (anonymous):

because it had a 45 mile head start

OpenStudy (anonymous):

then what is x and y

OpenStudy (anonymous):

i will let you figure that out you know the speed, you know the time, therefore you know the distance

OpenStudy (anonymous):

d=v*t

OpenStudy (anonymous):

what is the time

OpenStudy (anonymous):

3 pm according to the question